OpenStudy (anonymous):

What is the slope of the line 2x – 5y = 12 ? A. two fifths B. five halves C. 2 D. negative twelve fifths

OpenStudy (shayaan_mustafa):

do you know what is the general form of slope formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (shayaan_mustafa):

it is y=mx+b where, b=constant and m=slope of the function and y=function and x=variable.

OpenStudy (anonymous):

2/5

OpenStudy (shayaan_mustafa):

now make your given function i.e. 2x-5y=12 according to the given form, separate y, then we got it as, y=(2/5)x-12/5 now compare with the given form, m=2/5, b=12/5 That's it. Regards Electronics engineer.